# ]Meta GenAI

### v1

@@ Abstract
Current pre-trained large language models typically need instruction tuning to align with human preferences. However, instruction tuning data is often quantity-saturated due to the large volume of data collection and fast model iteration, leaving coreset data selection important but underexplored. On the other hand, existing quality-driven data selection methods such as LIMA (NeurIPS 2023 (Zhou et al., 2024)) and AlpaGasus (ICLR 2024 (Chen et al., )) generally ignore the equal importance of data diversity and complexity. In this work, we aim to design a diversity-aware data selection strategy and creatively propose using sparse autoencoders to tackle the challenge of data diversity measure. In addition, sparse autoencoders can also provide more interpretability of model behavior and explain, e.g., the surprising effectiveness of selecting the longest response (ICML 2024 (Zhao et al., )). Using effective data selection, we experimentally prove that models trained on our selected data can outperform other methods in terms of model capabilities, reduce training cost, and potentially gain more control over model behaviors.
@@ 1 Introduction
Scaling large language models (LLMs) has been shown to significantly enhance performance (Kaplan et al., 2020; Achiam et al., 2023), augmented by alignment (Bai et al., 2022; Rafailov et al., 2024; Schulman et al., 2017) to make LLMs follow human instructions. Instruction fine-tuning (IFT) (Wei et al., ; Longpre et al., 2023; Sanh et al., ; Ouyang et al., 2022) has become an essential step in adapting LLMs to perform effectively across diverse tasks, and it is believed that the data quality and diversity are the most important factors during IFT. Recently, there is a growing interest on data-centric AI. For example, innovations in data engineering enable scaling to vast contexts, extending the model’s capacity to handle extensive data sequences (Fu et al., ). Alpaca (Taori et al., 2023b) can elicit Llama’s instruction-following by distilling $52$k instruction conversations from ChatGPT (OpenAI, 2023). Techniques like importance resampling for data selection have been developed to optimize training data relevance (Xie et al., 2023).
@@
Previous work has proved the effectiveness of small data quantity for achieving either general instruction following (Chen et al., 2023; Zhou et al., 2024), or task-oriented abilities (Xia et al., ). Zhang et al. (2024a) shows some preliminarily results to claim that instruction diversity drives generalization to unseen scenarios through symbolic tasks. And Liu et al. perform controlled studies to measure data across three dimensions: complexity, quality, and diversity (measured by cosine similarity of sentence representation). Besides, data pruning based on concept complexity has been shown effective in both pre-training (Abbas et al., ) and post-training (Lu et al., ). However, many previous text diversity metrics (Shaib et al., 2024) are not adopted due to their intrinsic limitations. Therefore, there is still a lack of reliable diversity measure toward more efficient, effective, and targeted data selection in the instruction tuning landscape. And it is also unclear about the underlying mechanism such as what contributes to an accurate data diversity measure. Recently, sparse autoencoder (SAE) (Cunningham et al., 2023) has emerged as a powerful tool for interpreting the activated features in language models. The sparsity of the atomic monosemanticity guarantees the features are highly independent and accurate. Inspired by this advance, we first train a SAE and then propose a new approach to use the activated features from SAE to measure data diversity. Depending on the number of selected data, we then proposed two algorithms for effective data selection. The first one is greedy sampling using features from SAE for limited data (SAE-GreedSelect) and the second one is similarity-based sampling using features from SAE for scaling up the data selection (SAE-SimScale). We conduct comprehensive experiments on the widely used Alpaca (Taori et al., 2023a) and WizardLM_evol_instruct_70k (Xu et al., 2023) datasets and our approach witnesses significant superiority over several previous results such as #InsTag (Lu et al., ), Longest (Zhao et al., ) and Repr Filter (Liu et al., ) across various models and data scale. In addition, our approach can also explain why some of the previous method works by looking at the extracted features in the selected data.
In summary, we propose a novel method to use sparse autoencoders for feature extraction, and offer new algorithms for diversity-driven data selection, contributing to more effective and interpretable instruction tuning methodologies.
@@ 2 Related Work
Sparse Autoencoders have been crucial in advancing our understanding of neural representations (Gao et al., 2024; Paulo et al., 2024; Braun et al., 2024), as well as in scaling and enhancing the interpretability of LLMs (Cunningham et al., 2023; Foote et al., 2023). The foundational work in sparse coding, also known as dictionary learning, dates back to an overcomplete basis set proposed around 30 years ago (Olshausen and Field, 1997). Building on this, K-SVD (Aharon et al., 2006) was developed as an algorithm for designing overcomplete dictionaries, facilitating sparse representations in signal processing. Furthermore, K-sparse autoencoders (Makhzani and Frey, 2013) introduced a K-sparse constraint to enforce sparsity, leading to more accurate data representations. Bau et al. (2020) highlight the importance of understanding unit-specific functions for model interpretability. Additionally, Tonolini et al. (2020) propose that the visual cortex may employ such strategies to efficiently represent natural images. Recently, sparse dictionary learning has been applied to visualize transformer models (Elhage et al., 2022; Henighan et al., 2023), revealing that contextualized embeddings can be expressed as linear combinations of transformer factors (Yun et al., 2021). Research has also investigated polysemanticity and capacity in neural networks, finding that neurons often encode multiple concepts, which poses challenges for interpretability (Scherlis et al., 2022; Lieberum et al., 2024). Moreover, engineering monosemanticity has been explored to design neurons that represent single concepts, thereby enhancing interpretability (Jermyn et al., 2022). A structured mathematical framework has been proposed for analyzing transformer circuits (Elhage et al., 2021). It has been demonstrated that LLMs can elucidate the functions of neurons within themselves, offering a novel perspective on model introspection (Bills et al., 2023). Sparse probing techniques have identified significant neurons, providing case studies that underscore the utility of sparsity in interpretability (Gurnee et al., ). Improvements in dictionary learning with Gated-SAE (Rajamanoharan et al., 2024a) and JumpReLU-SAE (Rajamanoharan et al., 2024b) have further enhanced the quality of learned representations.
@@
Data Selection (Albalak et al., 2024; Wang et al., 2024), the process of selecting a representative subset of data to achieve efficient training without compromising model performance, is important in both pre-training (Brandfonbrener et al., 2024; Tirumala et al., 2023) and post-training (Chen et al., ; Li et al., 2024b). Previous studies (Mindermann et al., 2022; Paul et al., 2021) have focused on optimizing this selection for various training objectives, specifically targeting model performance constraints (Xia et al., 2024), the value of individual data points (Covert et al., ), and addressing bias (Jain et al., 2024). Recently, the emphasis on data curation (Taori et al., 2023b; Chiang et al., 2023; Cui et al., 2024; Wang et al., 2023) and selection (Zhou et al., 2024) for LLMs suggests that the main capabilities of LLMs come from pre-training, and a small amount of well-crafted instruction data can enable excellent instruction following. As a result, various data selection methods have been proposed. For example, AlpaGasus (Chen et al., ) uses ChatGPT to score data quality and selects only the top 1000 highest-scoring data points. Zhao et al. propose a simple yet effective baseline of selecting the longest responses, while Xia et al. ; Zhang et al. (2024b); Pan et al. (2024) employ gradient-based clustering for task-agnostic coreset selection. Further research (Liu et al., ) indicates that data quality (Li et al., 2024c; Ding et al., 2023; Li et al., 2024a, 2023), diversity (Ge et al., 2024), and complexity (Xu et al., 2023; Sun et al., 2024; Ivison et al., 2023) are all crucial to the success of IFT, especially under complex constraints. However, accurately measuring these dimensions is a nontrivial task. In this work, we focus on data diversity. Data diversity in instruction-tuning is increasingly recognized as crucial for building robust models (Bukharin and Zhao, 2023). #InsTag (Lu et al., ) measures data diversity through intention tags of instructions, a method also adopted in the official technical reports of advanced LLMs, including Llama-3 (Dubey et al., 2024) and Qwen2-72B (Bai et al., 2023; Yang et al., 2024). In this work, we prove that SAEs offer a better diversity measure.
@@ 3 Sparse Autoencoder
4:     Set $\mathcal{T}_{s}^{B}\leftarrow\emptyset$
@@
6:        if $|\mathcal{T}_{s}^{B}\cap\mathcal{T}_{q}|>|\mathcal{T}_{s}^{B}|$ (if Greedy: SAE-GreedSelect ) or $|\mathcal{T}_{s}^{B}\cap\mathcal{T}_{q}|/|\mathcal{T}_{s}^{B}|>\theta$ (if Similarity-based: SAE-SimScale ) then
@@ 4.1 Diversity-driven Data Selection through Sparse Autoencoder
Based on the extracted features from our trained SAEs, we design two sets of data selection methods: 1) when we only want to select a limited number of data as presented in (Chen et al., ; Zhao et al., ), we propose greedy sampling using features from SAE for limited data (SAE-GreedSelect) to maximize the utilization of features, and 2) when we want to scale up the selected data rather than only picking a fixed number of data, we propose similarity-based sampling using features from SAE for scaling up the data selection (SAE-SimScale). For example, the original #InsTag (Lu et al., ) method uses a greedy search but can not scale to $5$k data since no data can bring new intention tags. The details of the two methods can be found in Algorithm 1. As we mentioned above that we actually use JumpRelu during inference, and we empirically set the threshold to $10$ across all experiments in our work. For method SAE-SimScale, there is one additional parameter of similarity ratio, and we set it to $0.8$ for all experiments. Thus, there are only $1$ and $2$ tunable parameters for SAE-GreedSelect and SAE-SimScale, respectively. This way, we can maximize the simplicity. Additional experiments regarding the threshold can be found in Section 5.7, and some even witness better results.

### v2

@@ Abstract
Instruction tuning data are often quantity-saturated due to the large volume of data collection and fast model iteration, leaving data selection important but underexplored. Existing quality-driven data selection methods, such as LIMA (NeurIPS 2023 (Zhou et al., 2024)) and AlpaGasus (ICLR 2024 (Chen et al., )) generally ignore the equal importance of data diversity and complexity. In this work, we aim to design a diversity-aware data selection strategy and creatively propose using sparse autoencoders (SAEs) to tackle the challenge of data diversity measure. In addition, SAEs can also provide more interpretability of model behavior and explain, e.g., the surprising effectiveness of selecting the longest response (ICML 2024 (Zhao et al., )). Using effective data selection, we experimentally prove that models trained on our selected data can outperform other methods in terms of model capabilities, reduce training cost, and potentially gain more control over model behaviors. We prove that SAEs can serve as a good alternative to diversity measure and design our method to be scalable for potential industrial large-scale pruning, and we will also release our trained SAEs for use by the broader community.
@@ 1 Introduction
Scaling large language models (LLMs) has been shown to significantly enhance performance (Kaplan et al., 2020; Achiam et al., 2023), augmented by alignment (Bai et al., 2022; Rafailov et al., 2024; Schulman et al., 2017) to make LLMs follow human instructions. Instruction fine-tuning (IFT) (Wei et al., ; Longpre et al., 2023; Sanh et al., ; Ouyang et al., 2022) has become an essential step in adapting LLMs to perform effectively across diverse tasks, and it is believed that the data quality and diversity are the most important factors during IFT. Recently, there is a growing interest on data-centric AI. For example, innovations in data engineering enable scaling to vast contexts, extending the model’s capacity to handle extensive data sequences (Fu et al., ). Alpaca (Taori et al., 2023b) can elicit Llama’s instruction-following by distilling $52$k instruction conversations from ChatGPT (OpenAI, 2023). Importance resampling for selection has been developed to optimize data relevance (Xie et al., 2023).
@@
Previous work has proved the effectiveness of small data quantity for achieving either general instruction following (Chen et al., 2023; Zhou et al., 2024), or task-oriented abilities (Xia et al., ). Zhang et al. (2024a) shows some preliminarily results to claim that instruction diversity drives generalization to unseen scenarios through symbolic tasks. And Liu et al. perform controlled studies to measure data across three dimensions: complexity, quality, and diversity (measured by cosine similarity of sentence representation). Besides, data pruning based on concept complexity has been shown effective in both pre-training (Abbas et al., ) and post-training (Lu et al., ). However, many previous text diversity metrics (Shaib et al., 2024) are not adopted due to their intrinsic limitations. Therefore, there is still a lack of reliable diversity measure toward more efficient, effective, and targeted data selection in the instruction tuning landscape. And it is also unclear about the underlying mechanism such as what contributes to an accurate data diversity measure. Recently, SAEs (Cunningham et al., 2023) has emerged as a powerful tool for interpreting the activated features in language models. The sparsity of the atomic monosemanticity guarantees the features are highly independent and accurate. Inspired by this advance, we first train an SAE and then propose a new approach to use the activated features from SAE to measure data diversity. Depending on the number of selected data, we then proposed two algorithms for effective data selection. The first one is greedy sampling using features from SAE for limited data (SAE-GreedSelect) and the second one is similarity-based sampling using features from SAE for scaling up the data selection (SAE-SimScale). We conduct comprehensive experiments on the widely used Alpaca (Taori et al., 2023a) and WizardLM_evol_instruct_70k (Xu et al., 2023) datasets and our approach witnesses significant superiority over several previous results such as #InsTag (Lu et al., ), Longest (Zhao et al., ) and Repr Filter (Liu et al., ) across various models and data scale. Our approach also explains why some previous methods work by looking at the extracted features in the selected data.
In summary, we propose a novel method to use SAEs for diversity measure, and design new algorithms for diversity-driven data selection, contributing to more effective and interpretable instruction tuning.
@@ 2 Related Work
Sparse Autoencoders (SAEs) are powerful for understanding neural representations (Gao et al., 2024; Paulo et al., 2024; Braun et al., 2024), as well as in scaling and enhancing the interpretability of LLMs (Cunningham et al., 2023; Foote et al., 2023). The foundational work in sparse coding, also known as dictionary learning, dates back to an overcomplete basis set proposed around 30 years ago (Olshausen and Field, 1997). Based on this, K-SVD (Aharon et al., 2006) was developed as an algorithm for designing overcomplete dictionaries, facilitating sparse representations in signal processing. Furthermore, K-SAEs (Makhzani and Frey, 2013) introduced a K-sparse constraint to enforce sparsity, leading to more accurate data representations. Bau et al. (2020) highlights the importance of understanding unit-specific functions for the interpretability of the model. Furthermore, Tonolini et al. (2020) proposes that the visual cortex can employ such strategies to efficiently represent natural images. Recently, sparse dictionary learning has been applied to visualize transformer models (Elhage et al., 2022; Henighan et al., 2023), revealing that contextualized embeddings can be expressed as linear combinations of transformer factors (Yun et al., 2021). Research has also investigated polysemanticity and capacity in neural networks, finding that neurons often encode multiple concepts, which poses challenges for interpretability (Scherlis et al., 2022; Lieberum et al., 2024). Moreover, engineering monosemanticity has been explored to design neurons that represent single concepts, thus enhancing interpretability (Jermyn et al., 2022). A structured mathematical framework has been proposed for analyzing transformer circuits (Elhage et al., 2021). It has been demonstrated that LLMs can elucidate the functions of neurons within themselves, offering a novel perspective on model introspection (Bills et al., 2023). Sparse probing techniques have identified significant neurons, providing case studies that underscore the utility of sparsity in interpretability (Gurnee et al., ). Improvements in dictionary learning with Gated-SAE (Rajamanoharan et al., 2024a) and JumpReLU-SAE (Rajamanoharan et al., 2024b) further enhanced the quality of learned representations.
@@
Data Selection (Albalak et al., 2024; Wang et al., 2024), the process of selecting a representative subset of data to achieve efficient training without compromising model performance, is important in both pre-training (Brandfonbrener et al., 2024; Tirumala et al., 2023) and post-training (Chen et al., ; Li et al., 2024b). Previous studies (Mindermann et al., 2022; Paul et al., 2021) have focused on optimizing this selection for various training objectives, specifically targeting model performance constraints (Xia et al., 2024b), the value of individual data points (Covert et al., ), and addressing bias (Jain et al., 2024). Recently, the emphasis on data curation (Taori et al., 2023b; Chiang et al., 2023; Cui et al., 2024; Wang et al., 2023) and selection (Zhou et al., 2024) for LLMs suggests that the main capabilities of LLMs come from pre-training, and a small amount of well-crafted instruction data can enable excellent instruction following. As a result, various data selection methods (Du et al., 2023; Chen et al., 2023; Xia et al., 2024a; Ge et al., 2024; Lee et al., 2024; Liu et al., 2024) have been proposed. For example, AlpaGasus (Chen et al., ) uses ChatGPT to score data quality and selects only the top 1000 highest-scoring data points. Zhao et al. proposes a simple yet effective baseline of selecting the longest responses, while Xia et al. ; Zhang et al. (2024b); Pan et al. (2024) employ gradient-based clustering for task-agnostic coreset selection. Further research (Liu et al., ) indicates that data quality (Li et al., 2024c; Ding et al., 2023; Li et al., 2024a, 2023), diversity (Ge et al., 2024), and complexity (Xu et al., 2023; Sun et al., 2024; Ivison et al., 2023) are all crucial to the success of IFT, especially under complex constraints. However, accurately measuring these dimensions is a nontrivial task. In this work, we focus on data diversity. Data diversity in instruction-tuning is increasingly recognized as crucial for building robust models (Bukharin and Zhao, 2023). #InsTag (Lu et al., ) measures data diversity through intention tags of instructions, a method also adopted in the official technical reports of advanced LLMs, including Llama-3 (Dubey et al., 2024) and Qwen2-72B (Bai et al., 2023; Yang et al., 2024). In this work, we prove that SAEs offer a better diversity measure.
@@ 3 Sparse Autoencoder
4:     Set $\mathcal{T}_{s}^{B}\leftarrow\emptyset$$\mathcal{T}$ is the set of activated features
@@
6:        if $|\mathcal{T}_{s}^{B}\cup\mathcal{T}_{q}|>|\mathcal{T}_{s}^{B}|$ (if Greedy: SAE-GreedSelect ) or $|\mathcal{T}_{s}^{B}\cap\mathcal{T}_{q}|/|\mathcal{T}_{s}^{B}|<\theta$ (if Similarity-based: SAE-SimScale ) then
@@ 4.1 Diversity-driven Data Selection through Sparse Autoencoder
We only focus on diversity for the selection to prove the effectiveness of SAEs against existing diversity measures. Based on the extracted features from our trained SAEs, we design two data selection methods: 1) when we only want to select a limited number of data as presented in (Chen et al., ; Zhao et al., ), we propose greedy sampling using features from SAE for limited data (SAE-GreedSelect) to maximize the utilization of features, and 2) when we want to scale up the selected data rather than only picking a fixed number of data, we propose similarity-based sampling using features from SAE for scaling up the data selection (SAE-SimScale). For example, the original #INSTAG (Lu et al., ) method uses a greedy search but can not scale to $5$k data since no data can bring new intention tags. The details of the two methods can be found in Algorithm 1. Compared with previous methods used in the industrial model training pipeline (Lu et al., ), our method is designed to be simple and scalable so that it can also be used for large-scale training by the industry.
As we mentioned above that we actually use JumpRelu during inference, and we empirically set the threshold to $10$ across all experiments in our work. For method SAE-SimScale, there is one additional parameter of similarity ratio, and we set it to $0.8$ for all experiments. Thus, there are only $1$ and $2$ tunable parameters for SAE-GreedSelect and SAE-SimScale, respectively. This way, we can maximize the simplicity. Additional experiments regarding the threshold can be found in Section 5.7, and some even witness better results.